OpenStudy (anonymous):

Hey! I did a bit but I need a couple more problems :/ I don't understand this one at all..

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

http://www.mathopenref.com/constparallel.html

OpenStudy (anonymous):

That's not the same thing. @jigglypuff314

OpenStudy (mathmale):

Hello, Nick, The illustration shows the y-axis, the point P to the right of the y-axis, and a line drawn through P. Here are a couple of ways of responding to this question: (1) Using P as your axis, rotate the line through P until it is exactly parallel to the y-axis. If you really wanted to be fussy, you could take a good ruler and use it to ensure that every point on the line thru P is the same distance from the y-axis. (2) You could measure how far P is from the y-axis in the horizontal direction (perpendicular to the y axis or parallel to the x-axis). Then mark another point Q, either farther from the x-axis or closer to it than P), and then reposition the existing line through P so that it goes through the new point Q as well.

OpenStudy (anonymous):

So the answer would be "Measure the distance from the intersection of the transversal to Point P".

OpenStudy (mathmale):

A lot depends on what your teacher's expectations are. Were I your teacher, I'd accept a verbal answer such as one of those I've typed above.

OpenStudy (anonymous):

@jigglypuff314 :) You're such a big help!

OpenStudy (mathmale):

Nick: if you're allowed (or encouraged) to use mathematical methods to answer questions such as these, I'd suggest you apply the Pythagorean Theorem. You have a triangle with the lengths of the legs known and are to find the length of the hypotenuse.

jigglypuff314 (jigglypuff314):

\[distance formula: \sqrt{(x _{2}-x _{1})^{2} + (y _{2}-y _{1})^{2}}\]given points (x1, y1) and (x2, y2)

OpenStudy (mathmale):

JigglyPuff: You really are a big help. Your formula is exactly right for this situation.

jigglypuff314 (jigglypuff314):

thanks :) it's better to go short and sweet, kids don't like reading paragraphs these days ;)

OpenStudy (anonymous):

5.83 :) ew. paragraphs *shivers*

jigglypuff314 (jigglypuff314):

lol and correct :)

OpenStudy (anonymous):

Sorry about being AFK im just trying to solve a bit on my own. I don't want to just ask for answers all the time :)

jigglypuff314 (jigglypuff314):

don't worry ;) you being able to do a lot of it by yourself is a good thing :)

OpenStudy (anonymous):

A triangle can be formed with side lengths 120,85,and 220 True or false?

jigglypuff314 (jigglypuff314):

120 + 85 = 205 120 - 85 = 35 220 is Not between 35 and 205

OpenStudy (anonymous):

Got it ;) Classify the triangle based on it's side lengths 5,12, and 13.

jigglypuff314 (jigglypuff314):

you can try... does 5^2 + 12^2 = 13^2? if it works, then it's a right triangle...

OpenStudy (anonymous):

Got it! :)

jigglypuff314 (jigglypuff314):

area of trapezoid: (1/2)h(b1+b2) in other words, average of bases times height.

OpenStudy (anonymous):

What...?

jigglypuff314 (jigglypuff314):

(11.75 + 18)(1/2)(6.12) =

OpenStudy (anonymous):

91.035 :) This question is different.

OpenStudy (anonymous):

Would it be 119?

jigglypuff314 (jigglypuff314):

not quite... add the area of the trapezoid to the area of the triangle the area of the triangle would be = (1/2)(12)(13.42) =

OpenStudy (anonymous):

80.52?

jigglypuff314 (jigglypuff314):

then add that to the answer you got before

OpenStudy (anonymous):

171.555

jigglypuff314 (jigglypuff314):

That's what I got :)

OpenStudy (anonymous):

:)

jigglypuff314 (jigglypuff314):

30x + 11y = 112 2x + 3y = 12

OpenStudy (anonymous):

I get y=0 but that is wrong :/

jigglypuff314 (jigglypuff314):

I solved this by the elimination method 30x + 11y = 112 2x + 3y = 12 (multiply through by 15) -> 30x + 45y = 180 - (30x + 11y = 112) -> 0 + 34y = 68 then divide both sides by 34 then plug what you get for y back into 2x + 3y = 12

OpenStudy (anonymous):

y=2?

OpenStudy (anonymous):

and x=3 :)

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

:D:D:D:DD

jigglypuff314 (jigglypuff314):

the diagonals bisect each other so to get one side it would be... √(1.5^2 + 2^2)

OpenStudy (anonymous):

10? :)))))))))))))

jigglypuff314 (jigglypuff314):

correct :) and DG = AB angle GDB = angle BAD DB = DB

OpenStudy (anonymous):

SAS

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (mathmale):

Hi, Nick! Glad you're making progress. How about posting no more than 1 or 2 problems in any one separate problem solving session?

OpenStudy (anonymous):

Got it jiggly

jigglypuff314 (jigglypuff314):

so we all get a chance of earning more medals! :D lol jk It's one of the site rules... I think it so that moderators can check up on things better and so in case reinforcements have to be called to help on a problem they wouldn't have to read through everything 0.o I had to do that once... not fun...

OpenStudy (anonymous):

Should we go to another one ? :/ or just stay?

jigglypuff314 (jigglypuff314):

how about make it a habit to just post each question separately ;)

OpenStudy (anonymous):

Ok ;3

jigglypuff314 (jigglypuff314):

height = √(20^2 - 12^2) = area = base times height

OpenStudy (anonymous):

Wrong one :3

jigglypuff314 (jigglypuff314):

I forget what it's called >,< like the one like if the triangles are congruent then its angles are congruent...

OpenStudy (anonymous):

SSA :3

jigglypuff314 (jigglypuff314):

whoops looks like I was thinking too far ahead :P not SSA though... SSA is not a true postulate. try QU = AD QD = AU DU = DU

OpenStudy (anonymous):

SSS

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

AAS

jigglypuff314 (jigglypuff314):

angle TAM = angle MAH MA = MA angle TMA = angle HMA

jigglypuff314 (jigglypuff314):

well I just got confuzeled... >,< idk this one (got cross-eyed)

OpenStudy (anonymous):

A. Not Congruent B.SSA C.HL D.SAS

jigglypuff314 (jigglypuff314):

A!!!

OpenStudy (anonymous):

Wait really

jigglypuff314 (jigglypuff314):

\[\Huge\bf\color{#FF0000}{A}\]yeah it was why I was confused, cuz I was like... "but... but... they're not congruent D,x"

OpenStudy (anonymous):

hahaha xD

jigglypuff314 (jigglypuff314):

area = length times width perimeter = 2L + 2W 38 = 2(10) + 2W solve for W and use it in area formula

OpenStudy (anonymous):

90 :D

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

I'll do a bit on my own :) after this one :3

jigglypuff314 (jigglypuff314):

c = √ (5√2)^2 + (5√2)^2 )

OpenStudy (anonymous):

Idk what that is :O

jigglypuff314 (jigglypuff314):

(5√2)^2 = 50 50 + 50 = 100 c = √100 =

OpenStudy (anonymous):

10.. I still don't understand that one :/

jigglypuff314 (jigglypuff314):

it was the triangle rule things... so for the question... x = 14/2 √3